OpenStudy (anonymous):

solve the equation, sinx=sqrt(3) cosx, where 0<=x<=90

OpenStudy (anonymous):

tanx=sqrt3

hartnn (hartnn):

i was typing that ^

hartnn (hartnn):

so, x=... ?

OpenStudy (anonymous):

60?

hartnn (hartnn):

yes.

OpenStudy (anonymous):

ohh, why this easy question i need to ask -_-

hartnn (hartnn):

no problem :)

OpenStudy (anonymous):

For 0<=x<360, how many roots does the equation 3sin^(2)x+2sinx-1=0 have?

OpenStudy (anonymous):

3y^(2)+2y-1=0

OpenStudy (anonymous):

y=-1 or 1/3

hartnn (hartnn):

very similar to previous question... what you get ?

OpenStudy (anonymous):

x=-90..... 270?

hartnn (hartnn):

for sin x = -1 yes, x= 270 is correct. and you already know there will be 2 values of x for sin x = 1/3 right ?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

19.47

OpenStudy (anonymous):

90+19.47 =109.47 ??

hartnn (hartnn):

sin (109.47) is not 1/3

hartnn (hartnn):

\(\sin (180-x)=sin x\)

OpenStudy (anonymous):

180-19.47=160.53?

hartnn (hartnn):

yes.

OpenStudy (anonymous):

3?

hartnn (hartnn):

total 3, correct :)

hartnn (hartnn):

wait!

hartnn (hartnn):

0<=x<=90 -_-

hartnn (hartnn):

270 and 160 won't work only 1 , 19.

OpenStudy (anonymous):

hey, that's the first question

OpenStudy (anonymous):

is the second question: For 0<=x<360

hartnn (hartnn):

sorry! 3 is correct.

hartnn (hartnn):

thats why, new Q, new post, please ?

OpenStudy (anonymous):

ok